What quantity of 75 per cent acid solution must be mixed with a 30 per cent solution to produce 810 mL of a 50 per cent solution?

Answer by stanbon(75887)

What quantity of 75 per cent acid solution must be mixed with a 30 per cent solution to produce 810 mL of a 50 per cent solution?
----------------
acid + acid = acid
0.75x + 0.30(810-x) = 0.50*810
---
Multiply thru by 100 to get:
75x + 30*910 - 30x = 50*810
45x = 20*810
x = (4/9)(810)
x = 4*90
x = 360 mL (amt. of 75% solution needed for the mixture)
============================
Cheers,
Stan H.